Hi, I'm having trouble with this problem:
Graph the linear equation by finding the x and y intercepts.
y= 1/2x + 3/4
Thank you so much for your help!!

Answer by mananth(16946)

y= 1/2x + 3/4
Put x=0
we get y= 3/4
(0,3/4)
put y=0
we get
0= 1/2 x +3/4
1/2 x= -3/4
x= -3/2
(-3/2,0)